OpenStudy (anonymous):

What is the solution of the system? Solve using matrices. x+2y = -1 -x - 3y = -2

OpenStudy (anonymous):

{1 2 -1] [-1 -3 -2]

OpenStudy (anonymous):

did you type that out right?

OpenStudy (anonymous):

the answer choices are: [-7,3], [3,-7], [7,-3], and 'no solution'.

OpenStudy (anonymous):

pretty sure its no solution .

OpenStudy (anonymous):

plugging in those values won't work.

OpenStudy (anonymous):

okay. thank you very much.

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

-7 3 does.

OpenStudy (anonymous):

-7 + 2(3) = -1 and 7 -3(3) = -2/

OpenStudy (anonymous):

I just checked it myself. It's [-7,3], yes?

OpenStudy (anonymous):

[-7 3] yes, I'm sorry. apparently I'm off tonight.

OpenStudy (anonymous):

it's fine. can you help me with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve the system: 3x+y-4z= -30 3x+2y+2z= -8 5x+5y+z= -26

OpenStudy (anonymous):

the answers are: (-30,-8,-26) (4,-2,4) (-4,-2,4) and (-4,2,-4)

OpenStudy (anonymous):

the proper way to do these is to isolate and use smaller equations, but with them giving you answer choices you can just plug and check. so how about we both do it and you tell me what you get.

OpenStudy (anonymous):

okay then.

OpenStudy (anonymous):

are you going to be actually tested on these and do you want to learn how to properly do them without the answers?

OpenStudy (anonymous):

It's -4,-2, and 4.

OpenStudy (anonymous):

yeah thats what I got as well.